Get Help

Online: Request Help
Phone: (707) 826-HELP (4357)
Email: help@humboldt [dot] edu
In Person: Library 120 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
Other Aug 6 2013 - 13:09 Customers of SharePoint and Project Server The system has processes that are stuck in the queued and cannot be flushed manually.
Other Aug 1 2013 - 18:00 New functionality is being added to the myHumboldt portal as part of the Evolution 3 project. This includes the ability for all users to customize their home tab; new resources, business tools and appearance for faculty & staff users and new personalized academic progress and financial aid information for students.
CMS/Peoplesoft Jul 24 2013 - 18:00 All CFS Production (FCFSPRD) and CFS Data Warehouse application and database services users. During our regularly scheduled maintenance window, CMS will be installing updates to the Fees application.
Network Folders Jul 19 2013 - 17:00
Other Jul 21 2013 - 05:00 All Users Scheduled Maintenance
CMS/Peoplesoft Jul 17 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following item will be installed: Current SAP Unit Obligation Code.
CMS/Peoplesoft Jul 16 2013 - 09:00 All CSU PeopleSoft users Unknown, we will update when we have more information from CMS Central System appears to be back up, will add detail if we get it from CMS Central
Other Jul 11 2013 - 08:36 Campus Users A phishing attempt went out early this morning with a subject of "Important Message About Your Humboldt State University Account" and a "From" field of "Humboldt State University Account <websupport@humboldt.edu>" advising users to click a link named "https://www.humboldt.edu". Please have your users ignore and delete.
Campus Data Network Jul 10 2013 - 23:33 Occupants of Hadley House #56 ONLY House is connected via high speed DSL over aging copper lines, running at 15Mbps. DSL system is unable to sustain 15Mbps and had to be set to 10Mbps to restore connectivity. TNS is actively working on a fiber optic solution expected to go live in the next 6 months.
CMS/Peoplesoft Jul 10 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following items will be installed: GRE Test Load and a Mod to Grades HSU Processing.
Domain Name Service Jul 9 2013 - 17:00 All Users The Secondary DNS server will be migrated to a Virtual Server due to running on obsolete hardware. The actual cut-over is expected to take less than 5 minutes, followed by testing. The system has been moved, and tested.
Campus Data Network Jul 8 2013 - 14:00 Network Users in Hadley House only *** Systems were rebooted and Service has been restored. *** TNS is investigating. Check back for updates.
Other Jul 5 2013 - 16:45 Wireless and ResNet Users The ResNet/Wireless DHCP server will be migrated from running on obsolete hardware to a Virtual Machine. The DHCP Server has been migrated.
Campus Phone System Jul 3 2013 - 15:20 Some campus voicemail users TNS is still investigating. We will know more after updates run Thursday morning TNS is investigating. Check back for updates. *** Update Service was restarted. Problem has cleared. A scheduled upgrade of the application will be upcoming to resolve the issue.
CMS/Peoplesoft Jul 3 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following items will be installed: MP 9.0, HR retrofit, and Self Service retrofit.
Campus Data Network Jun 28 2013 - 09:21 Campus Users Attempting to Connect to Sites Off-Campus ** UPDATE Sunday 6/30 The main network link has been restored as well **** (12:31) redundant link is back up. Network connectivity restored. Check back for updates on the main campus network link ** Still No ETA -- Check back for updates ** We're in touch with CENIC currently and they are investigating. Check back for updates. We are back !!!
Off-Campus Internet Connection Jun 30 2013 - 23:00 Any network user trying to connect off campus Due to the Network Outage on Friday 6/28, maintenance needs to be performed. There should not be any affect to the campus, but there could be rolling outages during this time perioud. If you are working at this time, please save your work frequently.
Other Jun 27 2013 - 17:00 Students, Faculty, Staff Migration to new server. System is Available.
Other Jun 27 2013 - 17:00 All Users The aging hardware serving the Campus DHCP server (Except Housing) will be retired.
Google Apps - Mail Jun 26 2013 - 11:30 All Google Mail Users Help Desk confirmed this issue to be campus-wide. Opening a ticket with Google. Google opened Case #03233237, and confirmed that a lot of their customers were having the same problem. The problem has allegedly been fixed in the meantime.
Moodle Jul 21 2013 - 05:00 Moodle Users We will be installing a critical patch.
Moodle Jun 28 2013 - 20:00 ** THIS EVENT HAS BEEN CANCELLED ** Moodle Users ** THIS EVENT HAS BEEN CANCELLED -- WE WILL BE RESCHEDULING THIS MAINTENANCE ** Moodle will be down for patching from Friday 8:00pm to 8:00am. This patching is required to fix a critical bug.
Campus Data Network Jun 24 2013 - 13:15 Network Users at Natural History Museum only *** UPDATE Power has been restored in town and so the network service is back up and working *** -- TNS investigated and found there is a power outage in town. Unknown at this time how long the power will be out. -- Check back for updates.
Campus Data Network Jun 21 2013 - 12:53 Marine Lab network users only *** UPDATE - Service has been restored, TNS is still investigating the cause *** TNS is investigating. Check back for updates.
Campus Data Network Jun 19 2013 - 16:38 All Hadley House network users *** UPDATE - Service has been restored. *** TNS is investigating. Check back for updates.
CMS/Peoplesoft Jun 19 2013 - 18:00 All CFS Production (FCFSPRD) and CFS Data Warehouse application and database services users. During our regularly scheduled maintenance window the following items will be installed: PeopleSoft Fixes for 9.0 AP, AR Aging by Customer Report, Enterprise Fees, Instructional Equipment, Reporting Solution, and GAAP Reports.
Other Jun 18 2013 - 09:41 All Portal users HSU Portal server was temporarily not accessible for approximately 10 minutes. It is now back up and available.
Other Jun 17 2013 - 15:13 Sophos Home Users Sophos Home Users may have noticed in the past few days that there was a red "x" on the Sophos icon in their system trays indicating that Sophos could not contact the server. The system has been restarted and updating service is restored.
Other Jun 15 2013 - 13:00 All Nolij Users / OAK database server users Shutdown delays due to excessive Database server Memory Use.
Other Jun 15 2013 - 05:00 All Users Scheduled Maintenance
feedback